# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v2

@@ 1 Introduction
What are the goals in this article ? We develop our geometrical approach towards quantum phase transitions, addressing transport properties and responses to circularly polarized light. The quantum phase transition is induced through a fixed magnetic field $M$ polarized along $z$ direction. In previous works, we have shown how the topological invariant can also be re-written in terms of the spin responses at the poles related to a protected quantum dynamo effect [23, 4, 24], with applications on topological quantum wires [25, 26, 27]. In this work, we evaluate the spin response on the surface of the ball, which is similar to a Skyrmion for $M=0$ as shown in Fig. 1. We introduce the effective magnetic moment of the monopole as a function of $M$ associated to the spin response in Eq. (20). We show how this is related to the transport properties. We introduce the associated susceptibility for the monopole which remains invariant within the topological phase until the phase transition and therefore can be introduced as a topological marker. To the best of our knowledge, this magnetic susceptibility of the monopole was not mentioned before in the literature. The magnetic moment has an orbital interpretation in the sense of the angular momentum associated to the transverse pumped current (charge) and this topological susceptibility shows then an analogy with orbital magnetism where the quantum Hall response is also thought of as a topological orbital susceptibility [28] corresponding to the derivative of the orbital magnetic moment with respect to the applied perpendicular magnetic field. We will then develop the theory for the Haldane model in two dimensions [7] with the introduction of a Semenoff mass [29] which corresponds e.g. to the effect of a charge density wave substrate [30]. Topological properties describing the occurrence of an edge mode are generally robust until the phase transition. Above the transition, particles progressively reside on one site preferably as a result of the proximity effect referring to a trivial insulator which does not present an edge mode and is characterized through a zero index (invariant). Finding observables to describe such a transition towards this trivial phase is yet of great interest related to the relative populations on the two sublattices. We develop an analysis of observables in momentum and real space. We take time to present a numerical analysis showing how our geometrical approach from the poles and from the Dirac points is particularly efficient. We emphasize that Berry curvatures at the two Dirac points encode information on the quantum Hall conductivity locally in momentum space [6]. It is relevant to mention here recent progress on measuring locally topological or geometrical properties in momentum space from photo-luminescence related to the Bloch sphere approach [31].
Figure 1: Spin Response as a function of the external applied magnetic field $M$ on the Bloch sphere. When $M<B$, the sphere reveals the hedgehog structure related to a Skyrmion. At $M=B$ at the transition, properties of the spin response are mostly modified close to south pole where the polarization along z direction is effectively zero. When $M>B$ the spin polarization at south pole becomes identical to the spin polarization at north pole.
@@
Our geometrical approach can then characterize the presence of half topological invariants occurring in the model of two interacting Bloch spheres [24, 4]. This model can be realized in quantum circuits [19] and can be applied as a platform for quantum information through protected Majorana fermions [32]. The half topological invariant means that the surface radiating the Berry curvature is halved corresponding to an hemisphere, producing a half Skyrmion, and that the geometrical properties are locally resolved at one pole as a $\pi$ Berry phase or $\pi$ winding number [33]. For the situation of entangled Bloch spheres, this also builds a parallel with a situation of a half flux quantum in the sense of superconductivity compared to a metallic phase [33]. The half topological invariant also finds practical applications for quantum materials through the prediction of a quantum anomalous Hall semimetal [24, 34, 35, 36, 37] and a quantum spin Hall semimetal in two dimensions [38]. This also finds applications related to interaction effects in topological Kitaev p-wave superconducting wires [25, 26]. In this article, we elaborate on this correspondence for the quantum phase transition of the Haldane model with a charge density wave substrate. We develop the geometrical correspondence for coupled-planes materials, including the possibility of QSH effect [39], and show an application of the Ramanujan alternating infinite series as a $\frac{1}{2}$ topological invariant in real space [40].
@@ 2 Bloch Sphere as a Magnetic Monopole, Quantum Hall Response and Topological Magnetism
It should be emphasized that for classical magnetic monopoles produced e.g. by the same radial magnetic field, the magnetic field $M$ corresponds to an additional source to the Gauss theorem and produces an additional effective zero flux on a spherical surface i.e. the in-coming magnetic flux at south pole is compensated by the out-going magnetic flux at north pole. In that case, the $M$ term will not mediate a classical phase transition and will not modify the topological charge produced by the radial field.
@@
The Hamiltonian reads
$H=-{\bf B}\cdot\bm{\sigma},$ (1)
with the magnetic field of the form
${\bf B}=B\left(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta+\frac{M}{B}\right).$ (2)
Here, $\theta\in[0;\pi]$ and $\varphi\in[0;2\pi]$ are the polar and azimuthal angles respectively in spherical coordinates. The energy eigenvalues are
$E_{\pm}=\pm\sqrt{(B\cos\theta+M)^{2}+B^{2}\sin^{2}\theta}.$ (3)
The eigenstate with (lowest) energy $E_{-}=-|{\bf B}|$ takes the general form
@@
The Hilbert space is formed with the 2D eigenstates $|+\rangle$, $|$-$\rangle$ associated to the $z$ direction (or Pauli matrix $\sigma_{z}$) with spin eigenvalues +1+1 and 1-1 respectively. We emphasize that the topological invariant is generally thought of as the magnetic charge for monopoles through an analogy with the electromagnetism and with the electric field formed from a source (electric charge). In quantum physics the Berry curvature plays the role of the magnetic field and the equivalent to the vector potential is the local Berry gauge potential [5]. The integral of the Berry curvature on the surface can be equivalently introduced in spherical coordinates or through a cartesian metric from the simplification of the $\sin\theta$ function in the differential area and the $\frac{1}{\sin\theta}$ factor stemming in the definition of the Berry gauge potential component along the equatorial direction in spherical coordinates [19]. In this way, we introduce the Berry gauge potential as
@@
In principle, the general correspondence should be thought of as $B^{*}\cos\tilde{\theta}=B\cos\theta+M$ and we introduce $B^{*}\sin\tilde{\theta}=B\sin\theta$. To ensure the validity of the map and the correspondence of eigenenergies we introduce $B^{*}=-E_{-}(M)$ with $E_{-}\rightarrow 0$ at $\theta=\pi$ when $M=B$. On the other hand, the Berry gauge potential and the Berry curvature will not depend specifically on the precise form of $B^{*}$. Therefore, we can simply introduce the unit radial $\bf n$} vector for the geometrical correspondence
@@
Figure 2: (Left) Classical magnetic monopole. We generalize the classical analogue of $A_{\varphi}$ introduced in Eq. (10) of Ref. [4] as $A_{\varphi}=-\frac{\cos\theta}{2}-\frac{M}{4}\cos(2\theta)$ (for a unit sphere). The component induced from the $M$ term in Eq. (10) does not modify the topological invariant. (Right) Representation of the effective topological properties for the quantum situation when including the magnetic field $M$ through the (dressed) angle $\tilde{\theta}(\theta=\pi)$. For all the phases the north pole is identified with $\theta=0$ i.e. $\tilde{\theta}=0$. Within the topological phase $B<M$, the south pole corresponds to $\tilde{\theta}(\theta=\pi)=\pi$. For $M>B$, the north and south poles are identical and the equivalent geometry does not encircle the topological charge at the origin. For $M=B$, geometrical properties are equivalent to a half sphere corresponding to a half monopole or a half Skyrmion.
@@
which can be interpreted as a half Skyrmion. It is interesting to mention that $A_{\varphi}(\tilde{\theta}(\theta=0))=+\frac{1}{2}$ corresponds to a topological characterization of a $\pi$ Berry phase encircling the north pole [24, 33].
From Ehrenfest’s theorem $\langle\psi|\sigma_{z}|\psi\rangle=\langle\sigma_{z}\rangle=\cos\tilde{\theta}$. Then, the topological number can be reformulated through the spin polarizations at the north pole and at the dressed angle associated to $\theta=\pi$ :
@@
In this formulation, in the presence of the radial magnetic field acting on the spin of an electron and in the presence of the magnetic field source $M$, we measure the transverse pumped charge associated to an electric field applied along the polar angle such that $\hbar\dot{\tilde{\theta}}=eE$. At time $t=0$, a charge is present at north pole. The charge goes down similar to the apple and the gravitational force is a result of the electric field directed along the polar angle direction i.e. corresponds to a Coulomb force. The electron moves on the surface of this metallic ball associated to the Bloch sphere. The final time $T$ is fixed to be identical for any $M$ such that the angle $\tilde{\theta}$ reaches $\pi$ i.e. $T=\frac{eE}{\hbar\pi}=\frac{2eE}{h}$. Within the topological phase, the transverse pumped charge at south pole precisely measures the topological invariant itself $C=1$. When we reach the transition maintaining the same value of the (final) time $T$ for the measure, associated to Fig. 2, this corresponds to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=\frac{\pi}{2})=0$ for dressed angles $\tilde{\theta}\in[\frac{\pi}{2},\pi]$ such that there is no Berry curvature effectively present in the south hemisphere. In that case, the topological invariant becomes halved i.e. equal to $C_{1/2}$ and the transverse pumped charge is also halved. Above the transition, the charge remains at the north pole i.e. $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=0)$ and there is no transverse pumped current. In Appendix A, we also show the relation with the cylinder geometry.
In this article, then we introduce the mean of the transverse pumped charge on the original sphere associated to the polar angle $\theta\in[0;\pi]$
@@
When $M\rightarrow+\infty$, we also have that $\bar{Q}_{\perp}\rightarrow 0$ because $\langle\sigma_{z}\rangle=\cos\tilde{\theta}\rightarrow 1$. When $M\neq 0$, we introduce the formula
$\bar{Q}_{\perp}=\frac{eC(M=0)}{2}-\frac{\kappa(M)}{2},$ (19)
@@
Here, $\kappa$ is the mean of the spin response or the effective spin response as a function of $M$ associated to the surface area in Fig. 1. When $M=0$, from symmetry between the two hemispheres $\kappa=0$. In response to the radial magnetic field in Fig. 1 the two hemispheres produce an effective moment equal in amplitude but opposite in directions. We are precisely questioning the (topological) properties of $\kappa$ as a function of $M$ when inserting the equation
@@
$\langle\psi|\sigma_{z}|\psi\rangle=\cos\tilde{\theta}=\frac{B\cos\theta+M}{\sqrt{B^{2}\sin^{2}\theta+(B\cos\theta+M)^{2}}}.$ (21)
The linear response domain in $M$, i.e. assuming that $M\ll B$ within the topological phase, leads to
$\kappa=\frac{M}{B}\frac{2}{3}C=\frac{M}{B}\frac{2}{3},$ (22)
with the local form of the topological invariant in Eq. (13). At the transition point with $M=B$, we also find that
$\kappa(M=B)=\chi(M=B)=\frac{4}{3}C_{1/2}=\frac{2}{3},$ (23)
@@
with the half topological invariant introduced in Eq. (14). The linear response regime then remains applicable until the transition point. We verify this result with a numerical integration of $\frac{dI}{d\xi}$ with $I=2\kappa$ and $\xi=\frac{M}{B}$ in Fig. 3. A nice result of this article is then to introduce a topological magnetic susceptibility response for the hedgehog sphere (Skyrmion qubit) within the topological phase
$\chi=B\frac{\partial\kappa}{\partial M}=\frac{2}{3}C.$ (24)
When $M\rightarrow+\infty$ this results in $\langle\sigma_{z}\rangle\rightarrow 1$ such that $\kappa$ measures the unit area. The susceptibility will then tend to zero.
@@
Figure 3: Numerical evaluation of $\frac{dI}{d\xi}=2\chi$ as a function of $\xi$, that reproduces the analytical result $\frac{4}{3}C$ with $C=1$ and $\frac{8}{3}C_{1/2}$ with $C_{1/2}=\frac{1}{2}$ at the transition ; it remains perfectly quantized within the topological phase.

### v3

@@ 1 Introduction
What are the goals in this article ? We develop our geometrical approach towards quantum phase transitions, addressing transport properties and responses to circularly polarized light. The quantum phase transition is induced through a fixed magnetic field $M$ polarized along $z$ direction. In previous works, we have shown how the topological invariant can also be re-written in terms of the spin responses at the poles related to a protected quantum dynamo effect [23, 4, 24], with applications on topological quantum wires [25, 26, 27]. In this work, we evaluate the spin response on the whole surface of the ball, which is similar to a Skyrmion for $M=0$. We introduce the effective magnetic moment of the monopole as a function of $M$ associated to the spin response in Eq. (21). We show how this is related to the transport properties. This induces a susceptibility for the monopole in Eq. (33) which remains invariant within the topological phase until the phase transition and therefore can be introduced as a topological marker. To the best of our knowledge, this magnetic susceptibility of the monopole was not mentioned before in the literature. This topological susceptibility shows then an analogy with orbital magnetism where the quantum Hall response is also thought of as a topological orbital susceptibility [28] corresponding to the derivative of the orbital magnetic moment with respect to the applied perpendicular magnetic field. We will then develop the theory for the Haldane model in two dimensions [7] with the introduction of a Semenoff mass [29] which corresponds e.g. to the effect of a charge density wave substrate [30]. Topological properties describing the occurrence of an edge mode are generally robust until the phase transition. Above the transition, particles progressively reside on one site preferably as a result of the proximity effect referring to a trivial insulator which does not present an edge mode and is characterized through a zero index (invariant). Finding observables to describe such a transition towards this trivial phase is yet of great interest related to the relative populations on the two sublattices. We develop an analysis of observables in momentum and real space. We take time to present a numerical analysis showing how our geometrical approach from the poles and from the Dirac points is particularly efficient. We emphasize that Berry curvatures at the two Dirac points also encode information on the quantum Hall conductivity locally in momentum space within the parabolic band approximation model corresponding to the Dirac equation for the nearest-neighbor hopping term [6]. It is relevant to mention here recent progress on measuring locally topological or geometrical properties in momentum space from photo-luminescence related to the Bloch sphere approach [31].
@@
Our geometrical approach can then characterize the presence of half topological invariants occurring in the model of two interacting Bloch spheres or two interacting spins-1/2 [24, 4]. This model can be realized in quantum circuits [19] and can be applied as a platform for quantum information through protected Majorana fermions [32]. The half topological invariant means that the surface radiating the Berry curvature is halved producing a half Skyrmion and that the geometrical properties are locally resolved at one pole as a $\pi$ Berry phase or $\pi$ winding number [33]. For the situation of entangled Bloch spheres, this also builds a parallel with a situation of a half flux quantum in the sense of superconductivity compared to a metallic phase [33]. The half topological invariant also finds practical applications for quantum materials through the prediction of a quantum anomalous Hall semimetal [24, 34, 35, 36, 37] and a quantum spin Hall semimetal in two dimensions [38]. This also finds applications related to interaction effects in topological Kitaev p-wave superconducting wires [25, 26]. In this article, we elaborate on this correspondence for the quantum phase transition in the Haldane model with a charge density wave substrate. We develop the geometrical correspondence for coupled-planes materials, including the possibility of QSH effect [39], and show an application of the Ramanujan alternating infinite series as a physical $\frac{1}{2}$ topological invariant in real space [40].
@@ 2 Bloch Sphere as a Magnetic Monopole, Quantum Hall Response and Topological Magnetism
It should be emphasized that for classical magnetic monopoles produced e.g. by the same radial magnetic field, the magnetic field $M$ corresponds to an additional source to the Gauss theorem that produces an additional effective zero flux on a spherical surface i.e. the in-coming magnetic flux at south pole is compensated by the out-going magnetic flux at north pole. In that case, the $M$ term will not mediate a classical phase transition.
@@
We begin with the quantum Hamiltonian for the spin-1/2
$H=-(B{\bf e}_{r}+M{\bf e}_{z})\cdot\bm{\sigma}.$ (1)
Here, the vector $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is associated to Pauli matrices. The total magnetic field corresponds to the addition of a radial magnetic field $\bf B$} and of a magnetic field $M$ polarized along $z$ direction. Then, ${\bf e}_{r}$ and ${\bf e}_{z}$ are two unit vectors associated to these directions. The Hamiltonian can be equivalently written in spherical coordinates as
$H=-{\bf B}^{*}\cdot\bm{\sigma}$ (2)
with the magnetic field
${\bf B}^{*}=B\left(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta+\frac{M}{B}\right).$ (3)
Here, $\theta\in[0;\pi]$ and $\varphi\in[0;2\pi]$ are the polar and azimuthal angles respectively in spherical coordinates. In Fig. 1, we present the local variation of the unit vector ${\bf n}=\frac{{\bf B}^{*}}{B^{*}}$ as a function of $M$ on the unit sphere associated to the angles $\theta$ and $\varphi$. The energy eigenvalues are
$E_{\pm}=\pm\sqrt{(B\cos\theta+M)^{2}+B^{2}\sin^{2}\theta}=\pm B^{*}=\pm|{\bf B}^{*}|.$ (4)
The eigenstate with (lowest) energy $E_{-}=-B^{*}$ takes the form
@@
The Hilbert space is formed with the 2D eigenstates $|+\rangle$, $|$-$\rangle$ associated to the $z$ direction (or Pauli matrix $\sigma_{z}$) with spin eigenvalues +1+1 and 1-1 respectively. The eigenstates are formulated within a particular gauge choice for the phase related to the azimuthal angle. The theory presented below will be gauge invariant.
Figure 1: Local variation of the ${\bf n}=\frac{{\bf B}^{*}}{B^{*}}$ vector associated to the total magnetic field as a function of $M$ on the unit sphere associated to the angles $\theta$ and $\varphi$. Phase transition refers here to the situation of the topological quantum phase transition occurring when $M=B$ i.e. it will be associated to the jump of the topological invariant from one to zero. Associated to this transition, the $n_{z}$-component of this unit vector flips its direction at south pole.
In quantum physics the Berry curvature plays the role of the magnetic field and the equivalent to the vector potential is the local Berry gauge potential [5]. The topological invariant is measured from the integral of the Berry curvature on the surface. It equivalently measures a topological charge in the core of the sphere. It can be equivalently introduced in spherical coordinates or through a cartesian metric from the simplification of the $\sin\theta$ function in the differential area and the $\frac{1}{\sin\theta}$ factor stemming in the definition of the Berry gauge potential component along the equatorial direction in spherical coordinates [19]. In this way, we introduce the Berry gauge potential as
@@
In principle, the general correspondence should be thought of as $B^{*}\cos\tilde{\theta}=B\cos\theta+M$ and we introduce $B^{*}\sin\tilde{\theta}=B\sin\theta$. On the sphere described through the dressed angle $\tilde{\theta}$ and $\varphi$, the magnetic field is radial ${\bf B}^{*}=B^{*}{\bf e}_{r}$. The Berry gauge potential and the Berry curvature will not depend specifically on the precise form of $B^{*}$. Therefore, we equivalently introduce the unit radial $\bf n$$^{*}$ vector for this geometrical correspondence in Fig. 2(Right)
@@
Figure 2: (Left) Representation of the total magnetic field $\bf B$$^{*}$ on the unit sphere described through the angles $\theta$ and $\varphi$. (Right) Representation of the effective topological properties through the (dressed) angle $\tilde{\theta}(\theta)$. The direction of the total magnetic field $\bf B$$^{*}$ is then radial for all the values of $M$ associated to the unit vector $\bf{n}^{*}$ in Eq. (10). We introduce $\mathcal{A}=A_{\varphi}$ in Eq. (12). For all the phases the north pole is identified with $\theta=0$ i.e. $\tilde{\theta}=0$. Within the topological phase $B<M$, the south pole corresponds to $\tilde{\theta}(\theta=\pi)=\pi$. For $M<B$ topological properties are equivalent to those of a Skyrmion. For $M>B$, the north and south poles are identical and the equivalent geometry does not encircle the topological charge at the origin. For $M=B$, geometrical properties are equivalent to a half sphere corresponding to a half monopole or a half Skyrmion.
@@
which can be interpreted as a half Skyrmion. It is interesting to mention that $A_{\varphi}(\tilde{\theta}(\theta=0))=-\frac{1}{2}$ corresponds to a topological characterization of a $\pi$ Berry phase encircling the north pole [24, 33]. The classical analogue of the local Berry gauge potential i.e. the classical vector potential reads ${\mathcal{A}}=A_{\varphi}(M)=-B\frac{\cos\theta}{2}-\frac{M}{4}\cos(2\theta)$ when setting $r=1$. This generalizes the formula introduced in Eq. (10) in Ref. [4] for $M=0$. We verify that classically the induced $M$-component $\frac{M}{4}\cos(2\theta)$ remains identical at the two poles justifying why the Gauss theorem only measures the presence of the radial magnetic field for this situation.
From Ehrenfest’s theorem $\langle\psi|\sigma_{z}|\psi\rangle=\langle\sigma_{z}\rangle=\cos\tilde{\theta}=-2A_{\varphi}$. Then, the topological number can be reformulated through the spin polarizations at the north pole and at the dressed angle associated to $\theta=\pi$ :
@@
In this formulation, in the presence of the radial magnetic field acting on the spin of an electron and in the presence of the magnetic field source $M$, we measure the transverse pumped charge associated to an electric field applied along the polar angle such that $\hbar\dot{\tilde{\theta}}=eE$. At time $t=0$, a charge is present at north pole. The charge goes down similar to the apple and the gravitational force is a result of the electric field directed along the polar angle direction i.e. it corresponds to a Coulomb force. The electron moves on the surface of this ball associated to the Bloch sphere. The final time $T$, corresponding to the time of the measure, is fixed to be identical for any $M$ such that the angle $\tilde{\theta}$ reaches $\pi$ i.e. $T=\frac{eE}{\hbar\pi}=\frac{2eE}{h}$. Within the topological phase, the transverse pumped charge at south pole precisely measures the topological invariant itself $C=1$. When we reach the transition maintaining the same value of the (final) time $T$ for the measure, associated to Fig. 2, this corresponds to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=\frac{\pi}{2})=0$ for dressed angles $\tilde{\theta}\in[\frac{\pi}{2},\pi]$ such that there is no Berry curvature effectively present in the south hemisphere. In that case, the topological invariant becomes halved i.e. equal to $C_{1/2}$ and the transverse pumped charge is also halved. Above the transition, the charge remains at the north pole i.e. $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=0)$ and there is no transverse pumped current. In Appendix A, we also show the relation with the cylinder geometry.
In this article, then we introduce the mean value of the transverse pumped charge on the original sphere associated to the polar angle $\theta\in[0;\pi]$
@@
When $M\neq 0$, we introduce the formula
$\bar{Q}_{\perp}=\frac{eC(M=0)}{2}-\frac{e\kappa(M)}{2},$ (20)
@@
Here, $\kappa$ is the mean value of the spin response along $z$ direction or the effective spin response as a function of $M$. When $M=0$, from symmetry between the two hemispheres $\kappa=0$. In response to the radial magnetic field in Fig. 1 the two hemispheres produce an effective moment equal in amplitude but opposite in directions. When $M\rightarrow+\infty$, $\langle\sigma_{z}(\tilde{\theta})\rangle\rightarrow 1$ such that the mean value of the transverse pumped charge on the whole area is zero. In that limit, $\kappa$ is a measure of the unit area. We are precisely questioning the (topological) properties of $\kappa$ as a function of $M$ when inserting the equations
@@
$\langle\psi|\sigma_{z}|\psi\rangle=\cos\tilde{\theta}=\frac{B\cos\theta+M}{\sqrt{B^{2}\sin^{2}\theta+(B\cos\theta+M)^{2}}}=\frac{\partial B^{*}}{\partial M}=-\frac{\partial E_{-}}{\partial M}.$ (22)
To evaluate $\kappa(M)$, we will first write down a correspondence between spin response and magnetic field. We introduce the definition
$B^{*}=B\sqrt{1+x}$ (23)
with
$x=\left(2\frac{M}{B}\cos\theta+\frac{M^{2}}{B^{2}}\right).$ (24)
At fixed $M$, $2Md(\cos\theta)=Bdx$. In this way,
$\kappa(M)=-\frac{B}{2}\frac{\partial}{\partial M}\left(\frac{1}{2M}\int_{x_{min}=x(\theta=0)}^{x_{max}=x(\theta=\pi)}B^{*}dx\right).$ (25)
This is equivalent to
$\kappa(M)=-\frac{B^{2}}{2}\frac{\partial}{\partial M}\left(\frac{1}{2M}\int_{x_{min}=x(\theta=0)}^{x_{max}=x(\theta=\pi)}\sqrt{1+x}\ dx\right).$ (26)
Then, we can relate $x_{min}=2\frac{M}{B}+\frac{M^{2}}{B^{2}}$ and $x_{max}=-2\frac{M}{B}+\frac{M^{2}}{B^{2}}$ revealing the properties of the magnetic field $\bf{B}^{*}$ at the poles in Fig. 1 :
$\sqrt{1+x_{max}}=\frac{B^{*}(\theta=\pi)}{B}=\left|1-\frac{M}{B}\right|$ (27)
$\sqrt{1+x_{min}}=\frac{B^{*}(\theta=0)}{B}=\left(1+\frac{M}{B}\right).$ (28)
This leads to
$\kappa(M)=-\frac{B^{2}}{2}\frac{\partial}{\partial M}\left(\frac{1}{3M}\left|1-\frac{M}{B}\right|^{3}-\frac{1}{3M}\left(1+\frac{M}{B}\right)^{3}\right).$ (29)
For $M<B$, the only terms which are relevant are the quadratic terms in the parenthesis which then turn into a linear term when applying the partial derivative with respect to $M$. For $M<B$, we find
$\kappa(M)=\frac{M}{B}\frac{2}{3}.$ (30)
@@
When $M<B$, this results in a quantized plateau for the susceptibility response. The linear response regime then remains applicable until the transition point. We verify this result with a numerical integration of $\frac{dI}{d\xi}$ with $I=2\kappa$ and $\xi=\frac{M}{B}$ in Fig. 3. A nice result of this article is then to introduce a topological magnetic susceptibility response for the hedgehog sphere (Skyrmion qubit) within the topological phase
$\chi=B\frac{\partial\kappa}{\partial M}=\frac{2}{3}.$ (31)
The susceptibility has indeed a topological origin when $M<B$. To show this we can re-phrase the result when $M<B$ as
$\kappa=-\frac{1}{6B}\frac{\partial}{\partial M}\left(\left(1-\frac{M}{B}\right)M\right)+\frac{1}{6B}\frac{\partial}{\partial M}\left(\left(1+\frac{M}{B}\right)M\right).$ (32)
This is then equivalent to
$\chi=\chi_{topo}=B\frac{\partial\kappa}{\partial M}=\frac{2}{3}C(M<B)=-2B\frac{\partial\bar{Q}_{\perp}}{\partial M}.$ (33)
where the topological invariant is the invariant in the topological phase which can be written as
$C(M<B)=\frac{1}{2}(\langle\sigma_{z}(0)\rangle-\langle\sigma_{z}(\pi)\rangle)=1$ (34)
from Eq. (14) and we insert the definition in Eq. (22) for the spin responses. Within the topological phase $M=B^{-}$ the topological character of $\kappa$ can be summarized as
$\kappa=\kappa_{topo}=-\frac{1}{6B}\left(B^{*}(\theta=\pi)-B^{*}(\theta=0)\right)+\frac{M}{3B}C(M<B).$ (35)
Since
$\frac{1}{2}\left(B^{*}(\theta=0)-B^{*}(\theta=\pi)\right)=MC(M<B)$ (36)
then, this also leads to
$\kappa=\kappa_{topo}=\frac{2M}{3B}C(M<B).$ (37)
This formula is valid until $M=B^{-}$. For $M>B$, the form of $\kappa$ in Eq. (29) is
$\kappa(M)=1-\frac{1}{3}\frac{B^{2}}{M^{2}},$ (38)
which leads to a power-law susceptibility response in $M^{-3}$. For $M=B$, the mean value of the magnetic moment and the susceptibility are continuous.
@@
Figure 3: Numerical evaluation of $\frac{dI}{d\xi}=2\chi$ as a function of $\xi$ that reproduces the analytical results, in particular the quantized plateau for the susceptibility within the topological phase.
Then, we present another view of this proof related to the properties of the topological invariant within the topological phase. We develop the $\sqrt{1+x}$ function in Eq. (26) as an infinite series and we develop the response order by order in $x=2\frac{M}{B}\cos\theta+\frac{M^{2}}{B^{2}}$. To order $\frac{x}{2}$, this leads to
$\kappa_{\frac{x}{2}}=\frac{M}{B}C(M=0)$ (39)
with the topological invariant
$C(M=0)=\frac{1}{2}\int_{0}^{\pi}\sin\theta d\theta=\frac{1}{2}\int_{0}^{\pi}\sin\tilde{\theta}d\tilde{\theta}=+1.$ (40)
The second equality reveals $C(M=0)=C(M<B)$ and imply that we study the topological phase. The linear term in $M$ in $\kappa$ coming from $\frac{x^{2}}{8}$ in the series associated to $\sqrt{1+x}$ then reads
$\kappa_{-\frac{x^{2}}{8}}=-\frac{1}{3}\frac{M}{B}C(M=0).$ (41)
Therefore,
$\kappa_{\frac{x}{2}}+\kappa_{-\frac{x^{2}}{8}}=\frac{2}{3}\frac{M}{B}C(M=0).$ (42)
Then, we verify that this result is in fact robust when including higher order terms in the series. This requires a systematic approach that the numerical method also reproduces. Therefore, we verify in this way that
$\kappa(M<B)=\kappa_{\frac{x}{2}}+\kappa_{-\frac{x^{2}}{8}}=\frac{2}{3}\frac{M}{B}C(M<B)=\frac{2}{3}\frac{M}{B}.$ (43)